OpenStudy (anonymous):

Solve. 3x – 5 = 7x – 17

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

3(3)-5=7(3)-17 9-5=21-17 4=4

OpenStudy (anonymous):

thus x=3

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

3x - 5 = 7x -17 -5+17 = 7x-3 12=4x divide it by 4 x=3